OpenStudy (anonymous):

A candy clerk has three times as many kilograms of chocolate creams as peanut clusters. The clerk has 48kilograms of the two candies together. How many kilograms of peanut clusters does the clerk have?

OpenStudy (texaschic101):

c + p = 48 --> c = 48 - p c = 3p now sub 48 - p in for c in the 2nd equation 48 - p = 3p ---add p to both sides 48 = 3p + p 48 = 4p -- divide by 4 12 = p c + p = 48 c + 12 = 48 c = 48 - 12 c = 36 The candy clerk has 12 kg of peanut clusters and 36 kg of choc. cream

OpenStudy (anonymous):

thank you texaschic101